OpenStudy (anonymous):

FEEL LIKE GIVING OUT MEDALS !! :)) 2.8 is what percent of 4?

OpenStudy (anonymous):

0.112

OpenStudy (anonymous):

(4/100) of Y = 2.8, where Y is the unknown number. To find Y, we simply re-adjust the expression to Y = 2.8 x 100/4. This leads to Y = (2.8/4) x 100 = 0.7 x 100 = 70. 70.

OpenStudy (anonymous):

That's wrong :/

OpenStudy (anonymous):

THANK YOU ❤️

OpenStudy (anonymous):

no problem.

OpenStudy (anonymous):

Folks, we're supposed to be helping each other, not providing answers for one another.

OpenStudy (anonymous):

@ospreytriple | I showed her how to do the work. Therefore I helped her.

OpenStudy (anonymous):

I like when people give answers with an explanation to it :)

OpenStudy (anonymous):

Folks, we're supposed to be helping each other, not providing answers for one another.

OpenStudy (anonymous):

The answer is 2.8 is 70% of 4